# Fast nonparametric feature selection with error control using integrated path stability selection

### v1

@@ 4.  Simulation study
Among the other methods with theoretical error control, IPSSRF performs well in terms of identifying true positives while controlling false positives, though not as well as IPSSGB. IPSSL, whose model assumptions are violated, performs poorly, neither controlling false positives at the target E(FP) or FDR levels, nor identifying many true ones. SSBoost, whose curves are obtained by varying the target E(FP) between 0 and 10, has few false positives but also few true positives. Observe, for example, that the for SSBoost in Figure 3 is always significantly lower than those for IPSSGB and IPSSRF, despite all three methods having the same target E(FP). This is due in part to the weakness of the efp scores used by SSBoost relative to those used by IPSS (Melikechi and Miller, 2024). KOBT performs poorly in general; most notably, its actual FDRs far exceed its target FDRs of 0.1 and 0.3 in all of the simulation settings in Figure 4. KOBT also runs in around 90 seconds, making it one of the slowest methods. With the exception of IPSSL for classification, all of the stability selection-based methods run in at most 40 seconds, often less.
The absence of error control for Boruta and Vita is clearly apparent in Figures 3 and 4. Both methods usually have FDRs over 0.75—often 20 or more false positives on average—far surpassing IPSSGB and IPSSRF. Despite this, Boruta and Vita identify fewer true positives than IPSSGB with a target E(FP) of 3 and, in many cases, IPSSGB with a target FDR of 0.3. The RF and XGBoost curves are obtained by increasing the number of features selected, as ranked by their importance scores, from 0 to 30; that is, the first nonzero point along the curve is obtained by keeping only the feature with the highest importance score, the second point is obtained by keeping the features with the two largest importance scores, and so on. Interestingly, RF and XGBoost significantly underperform IPSSGB and IPSSRF in terms of true and false positives across the whole 0 to 30 range, indicating that IPSS is able to identify important features that simple thresholding of the baseline algorithm cannot, no matter how many features are kept. In terms of runtime, RF and XGBoost run quickly except in the $p=5000$ regression experiment. Boruta and Vita are typically slower than the IPSS methods, and this disparity grows with the number of features.
Figure 3. TP versus FP for the six simulation experiments. Columns correspond to the number of features, $p$, and the top and bottom rows correspond to regression and classification, respectively. Curves for the three IPSS methods and SSBoost are obtained by varying the target E(FP), with showing results when the target E(FP) is 3. The KOBT curve is obtained by varying the target FDR, and the RF and XGBoost curves are obtained by increasing the number of features selected (as ranked by their importance scores) from 0 to 30. The remaining methods, indicated by the symbols, do not have curves because they select fixed sets of features.
Figure 4. TPR versus FDR for the six simulation experiments. Results are similar to those in Figure 3, except now we show the FDR and TPR rather than E(FP) and E(TP), and curves for the three IPSS methods and KOBT are obtained by varying the target FDR, with and showing results when the target FDR is 0.1 and 0.3.
Figure 5. Runtimes for the six experiments. Columns correspond to the number of features, $p$, and the top and bottom rows correspond to regression and classification, respectively. Orange lines denote median runtime and whiskers extend to 1.5 times the interquartile range.
5.  Ovarian cancer
We apply IPSSGB, IPSSRF, and the other methods in Table 1 to microRNA (miRNA) data from the cohort of ovarian cancer patients described in Section 4 (Vasaikar et al., 2018). We report two studies here; see Section S3 for results from additional studies, including analyses of the RNA-seq data from this cohort. In the first study, the features are the 588 miRNAs and the response variable is prognosis, which is defined as whether the patient was still alive at the time of last follow-up. In the second study, the response is one of the miRNAs, miR-150, identified as important in the first stage, and the features are the remaining miRNAs. Thus, since miRNA measurements are continuous, the first study is a binary classification problem, and the second study is regression.
@@ 5.  Ovarian cancer
All analyses, both here and in Section S3, are conducted using 10-fold cross-validation (CV) as follows. First, all patients with missing values are removed, resulting in $n=442$ and $n=453$ for the first and second studies, respectively. The remaining patients are evenly partitioned into 10 groups. In each of the 10 CV steps and for each selection method, one group of patients is set aside (the test set), and a set of miRNAs is selected by applying the method to the data in the remaining groups (the training set). For the IPSS methods, we also compute the efp scores and $q$-values of each miRNA on the training set. Next, for each method we construct three predictive models—a linear model (either lasso or $\ell^{1}$-regularized logistic regression), a random forest model, and a gradient boosting model—using only the miRNAs selected by that method. Each model is then used to predict responses from the test set, and the smallest of the three prediction errors is recorded. We implement all three models so that no method has an inherent advantage over another. For example, the miRNAs selected by IPSSL may be better suited to minimizing error in a linear model than those selected by IPSSGB, while the miRNAs selected by IPSSGB may be better suited to minimizing error in a gradient boosting model than those selected by IPSSL. The linear and random forest predictive models are implemented with scikit-learn (Pedregosa et al., 2011) and gradient boosting with XGBoost (Chen and Guestrin, 2016), always with default parameters. For continuous responses, in each CV step we subtract the mean of the training responses from all responses, training and test, and scale all responses by the empirical standard deviation of the training responses.
Table 2 shows the top five miRNAs related to prognosis for each IPSS method, as ranked by their median $q$-values over the 10 CV steps. In Section S3, we further investigate the findings in Table 2 via a literature review. Notably, we find numerous studies that link several of the top miRNAs selected by IPSSGB and IPSSRF, namely miR-30d, miR-93, miR-96 and miR-150, to ovarian cancer prognosis. Among these, IPSSL only ranks miR-96 in its top five, and we found little literature identifying its top-ranked feature, miR-1-2, to ovarian cancer. The results of the second study, identifying miRNAs that are related to miR-150, are shown in Figure 6. Here, IPSSGB achieves the lowest average prediction error over the 10 cross-validation steps of any method, while using just 15 to 20 miRNAs, achieving lower prediction error than even the method that uses all 587 miRNAs to predict miR-150 levels (the dashed gray line in Figure 6). Figure 7 is similar to Figure 6, except now the response is tumor purity, the proportion of cancerous cells in a tissue sample. Here, we see IPSSGB performs best with the fewest features, while the top 30 features selected by IPSSRF achieve essentially identical predictive performance to the method that uses all 588 miRNAs.
IPSSGB IPSSRF IPSSL
miR-96(0.2) miR-1270(0.27) miR-1-2(0.07)
miR-150(0.28) miR-93(0.36) miR-1270(0.08)
miR-1270(0.43) miR-30d(0.36) miR-96(0.09)
miR-30d(0.43) miR-342(0.36) miR-342(0.09)
miR-1301(0.46) miR-150(0.41) miR-934(0.12)
Table 2. MicroRNA and prognosis. The top five miRNAs related to prognosis according to each IPSS method, as ranked by their median $q$-values over the 10 cross-validation steps.
@@
Figure 6. MicroRNA and miR-150. The horizontal and vertical axes show the average number of features selected and the average prediction error over the 10 cross-validation steps, respectively. Curves for the IPSS methods and SSBoost are obtained by varying the target E(FP) between 1/8 and 10. Points for these methods correspond to target E(FP)s of 1/8, 1/4, 1/2, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, with larger target E(FP)s generally corresponding to more features selected and lower prediction error. Similarly, points for KOBT correspond to target FDRs of 0.01, 0.025, 0.05, 0.1, 0.15, 0.2, 0.25, 0.3, and 0.4. The points for RF and XGBoost correspond to selecting between 1 and 30 features (hence, the number of features selected by these two methods is deterministic). Boruta and Vita do not have error control parameters and are therefore indicated by single points rather than curves.
Figure 7. MicroRNA and tumor purity. This is the same as Figure 6, except now the response is tumor purity, the proportion of cancerous cells in a sample. We again see the IPSS methods outperform other methods, achieving similar or better predictive performance with relatively few features.
@@ 6.  Discussion
We have demonstrated that IPSSRF and IPSSGB achieve superior results in terms of false positive control, detecting true positives, and computation time. More broadly, IPSS for thresholding is a general framework whose theory and implementation apply to arbitrary importance scores. For instance, examples of other scores include $p$-values (with smaller $p$-values indicating greater importance), Shapley values (used to quantify the contribution of individual features to neural networks and other machine learning models), and loadings in principal components analysis (which quantify the contribution of each feature to a given principal component). The main practical limitation to consider is the cost of computing the relevant importance scores, since IPSS must compute these scores on many random subsamples of the data.
We have also introduced efp scores and shown that, in addition to controlling E(FP), they can control the FDR and estimate $q$-values. Storey (2003) showed that $q$-values admit a Bayesian interpretation, suggesting a link between IPSS and Bayesian feature selection that could be an interesting line of future work. It would also be interesting to study whether $q$-values estimated from efp scores are more accurate than $q$-values estimated from other quantities, such as $p$-values (Storey and Tibshirani, 2003).
Finally, a more ambitious goal is to extend IPSS to unsupervised feature selection problems (that is, feature selection when there is no response variable) and non-iid data. The PCA-based scores mentioned above provide at least one way to apply IPSS in an unsupervised setting. Developing a rigorous approach to IPSS for non-iid data could provide novel methods for nonparametric feature selection with error control for time series, networks, and longitudinal data.
Data availability and code
All of the code and data used in this work are freely available at https://github.com/omelikechi/ipss. A Python package for implementing integrated path stability selection, including IPSSGB and IPSSRF, is freely available at https://pypi.org/project/ipss/
@@ Acknowledgements
Supplementary material for “Fast nonparametric feature selection with error control using integrated path stability selection”
S1.  Additional IPSS details
We provide an algorithm that implements integrated path stability selection (IPSS) for thresholding (Section S1.1), elaborate on the theory of IPSS and its connection to efp scores (Section S1.2), and describe and study the IPSS parameters in greater detail (Section S1.3).
S1.1.  Algorithm.
Algorithm S1, discussed in Section 2.3, implements IPSS for thresholding. The number of grid points used to evaluate the integrals in Algorithm S1 is always $K=100$. Like many of the other IPSS parameters, $K$ is inconsequential provided it is sufficiently large; in our experience, values greater than 25 suffice. This is because the function $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$, the paths $\lambda\mapsto\hat{\pi}_{j}(\lambda)$ (which are monotonically increasing functions of $\lambda$), the quantity $\mathcal{I}(\Lambda)$, and the measure $\mu$ are all very numerically stable.
Algorithm S1 Integrated path stability selection for thresholding
1:Data $Z_{1:n}$, importance function $\Phi$, number of grid points $K$, and number of iterations $B$.
@@
10:while $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])<0.05$ do
@@
S1.2.  Theory: From IPSS to efp scores.
@@
S1.3.  Parameters.
As noted above, the function $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ used when implementing IPSS in the main text is determined by the availability and strength of the theoretical bound in Theorem S1.1. Similarly, the choice of $\lfloor n/2\rfloor$ samples used to construct the selection probabilities $\hat{\pi}_{j}(\lambda)$ are required for stability selection theorems (not just Theorem S1.1) to hold, and it is unclear how to adapt their proofs to accommodate other sample sizes. Thus, $f$ and $\lfloor n/2\rfloor$ are theoretical necessities rather than free parameters.

### v2

@@ 3.  Simulation studies
Among the other methods with theoretical error control, IPSSRF performs well in terms of identifying true positives while controlling false positives, though not as well as IPSSGB. IPSSL1, whose parametric assumptions are violated, performs poorly, failing to control false positives at target FDR levels. It also identifies essentially no true positives in the Gaussian experiments. SSBoost is overly conservative, undershooting the target FDR at the expense of identifying few true positives. This is partly due to the weakness of the efp scores used by SSBoost relative to those used by IPSS (Melikechi and Miller, 2024).
All versions of model-X knockoffs underperform IPSSGB and IPSSRF in all experiments. This is even true in the Gaussian setting, where the known joint distribution of the features was used to implement these methods (recall that model-X knockoffs requires knowledge of the joint distribution, while IPSS does not). With the exception of KOL and KORF in the $p=5000$ and, to a lesser extent, $p=2000$ RNA-seq experiments, all of these methods control the FDR at target levels. KORF is the only knockoffs-based method that identifies essentially any true positives in the Gaussian experiments. DeepPINK, which combines knockoffs with deep neural networks, has the lowest TPR out of every method in almost all experiments. This agrees with our earlier observation that deep learning typically underperforms tree-based methods on tabular data and requires tens of thousands of samples for competitive performance (Grinsztajn et al., 2022).
The absence of error control for Boruta and Vita is clearly apparent in Figures 2 and S2. Both methods usually have FDRs over 0.75, far surpassing other methods. Despite this, Boruta and Vita almost always identify fewer true positives than IPSSGB when the target FDR is greater than 0.2 or, in some cases, even 0.1. Boruta and Vita are also slower than other methods, and this disparity grows with the number of features (Tables S3 and S4).
VSURF usually has a lower FDR than Boruta and Vita, but still underperforms IPSSGB. Its excessive runtimes prevented us from including it in the $p=2000$ and $5000$ RNA-seq experiments. Several other methods are also omitted, namely KOBT, RFEGB, and RFHT. Briefly, KOBT far exceeded target FDRs despite extensive tuning, while RFEGB and RFHT had FDRs over $0.8$ and extremely long runtimes. For details, see Sections S3 and S2.
Figure 2. RNA-seq simulation results (regression). First, second, and third columns correspond to the $p=500$, $2000$, and $5000$ experiments, respectively. The horizontal axis in each plot shows the target FDR. Methods without false discovery control (gray horizontal lines) do not vary with the target FDR. The black dashed line in the FDR plots represents perfect FDR calibration, FDR = target FDR. Non-gray solid lines correspond to IPSS or stability selection-based methods. Non-gray dashed lines show methods based on model-X knockoffs.
4.  Cancer studies
We study ovarian cancer and glioma (a type of brain cancer). Both studies include multiple substudies in which the features are either genes, measured by RNA-seq, or microRNAs (miRNAs). The response variables are either prognosis (whether the patient was alive at last follow-up), tumor purity (the proportion of cancerous cells in a tissue sample), or the expression level of a particular gene or miRNA. All data are from The Cancer Genome Atlas (Weinstein et al., 2013) and were downloaded from LinkedOmics (Vasaikar et al., 2018). Additional details and results are in Section S4.
We assess feature selection performance by (i) performing literature searches and (ii) assessing predictive performance using cross-validation (see Sections S4.3 and S4.4, respectively). The literature search results show that IPSSGB and IPSSRF consistently identify more known important features at lower target FDRs than other methods. In Table S6, for example, IPSSGB and IPSSRF identify 8 and 6 miRNAs, respectively, all but one of which have been implicated in ovarian cancer prognosis. In contrast, IPSSL1 identifies 4 miRNAs, missing the three most significant ones, while KOGLM, KORF, KOL1, DeepPINK, and SSBoost select no miRNAs at all, even when the target FDR is $0.5$.
In the RNA-seq and glioma prognosis study (Table S10), only IPSSGB identifies the key oncogene FOXM1 (Raychaudhuri and Park, 2011), and only IPSSGB, IPSSRF, and KORF identify WEE1, which is also known to play a significant role in glioma outcomes (Music et al., 2016). Furthermore, IPSSGB and IPSSRF are more confident in their selections, assigning WEE1 $q$-values of 0.10 and 0.06, respectively, while KORF does not select WEE1 until the target FDR is reduced to 0.44. More generally, Table S10 shows that IPSSGB, IPSSRF, and, to a lesser extent, IPSSL1, tend to identify genes supported by the glioma literature while avoiding genes with little or no known connection. In contrast, KORF and especially KOL1 select many genes with limited or no literature support, while KOGLM, DeepPINK, and SSBoost select no genes at all.
@@ 4.  Cancer studies
Briefly, our cross-validation (CV) studies proceed as follows. In each of 20 CV steps, one group of patients is set aside (the test set), and a set of features is selected by each method using the data in the remaining groups (the training set). Next, for each method we construct three predictive models—a linear model, a random forest model, and a gradient boosting model—using only the features selected by that method on the training data. Each model is then used to predict responses from the test set, and the smallest of the three prediction errors is recorded (we use mean squared error for regression and $1-\mathrm{accuracy}$ for classification). All three models are implemented so that no method has an inherent advantage over another. For example, features selected by IPSSL1 may be better suited to linear model predictions than those selected by IPSSGB, while those selected by IPSSGB may be better suited to gradient boosting predictions than the ones selected by IPSSL1.
Figure 3 shows the results from our RNA-seq and glioma prognosis CV study. On average, the top 20 genes selected by IPSSGB yield the same prediction error as the full model that uses all $10,058$ genes. IPSSRF and IPSSL1 achieve the smallest prediction errors overall, and only use 10 to 20 genes to do so. KORF and KOL1 have higher predictive errors despite using more than 40 selected genes, and Boruta, which selects over 100 genes, has an average error similar to that of the full model. Vita (not shown) selects over 500 genes on average and has an average prediction error of 0.22, and DeepPINK and SSBoost select no genes at all, even when the target FDR is 0.5.
@@
Figure 3. RNA-seq and glioma prognosis. The horizontal and vertical axes show the average number of genes selected and the average prediction error over the 20 cross-validation steps, respectively. Curves for each method are obtained by varying the target FDR between 0 and 0.5. DeepPINK and SSBoost selected no genes and are therefore shown by single points rather than curves. Boruta is also indicated by a single point because it does not have FDR control parameters. The dashed black line shows the average error when using all $p=10,058$ genes to predict prognosis.
@@ 5.  Discussion
We have demonstrated that IPSSRF and IPSSGB achieve superior results in terms of false positive control, true positive detection, and computation time. More broadly, IPSS for thresholding is a general framework whose theory and implementation apply to arbitrary importance scores. For instance, examples of other scores include $p$-values (with smaller $p$-values indicating greater importance), Shapley values (used to quantify the contribution of individual features to neural networks and other machine learning models), and loadings in principal components analysis (which quantify the contribution of each feature to a given principal component). The main practical limitation to consider is the cost of computing the relevant importance scores, since IPSS must compute these scores on multiple subsamples of the data.
We have also introduced efp scores and shown that, in addition to controlling E(FP), they can be used to control the FDR and estimate $q$-values. Storey (2003) showed that $q$-values admit a Bayesian interpretation, suggesting a link between IPSS and Bayesian feature selection that could be an interesting line of future work.
Finally, a more ambitious goal is to extend IPSS to unsupervised feature selection problems (that is, feature selection when there is no response variable) and non-iid data. The PCA-based scores mentioned above provide at least one way to apply IPSS in an unsupervised setting. Developing a rigorous approach to IPSS for non-iid data could provide novel methods for nonparametric feature selection with false discovery control for networks and spatially or temporally-indexed data.
Code and data
All code and data used in this work are available on GitHub (https://github.com/omelikechi/ipss_bioinformatics) and permanently archived on Zenodo (https://doi.org/10.5281/zenodo.15335289). A Python package for implementing IPSS is available on GitHub (https://github.com/omelikechi/ipss) and PyPI (https://pypi.org/project/ipss/). An R implementation of IPSS is also available on GitHub (https://github.com/omelikechi/ipssR).
@@ Acknowledgements
SUPPLEMENTARY MATERIALS
We provide further details about IPSS (Section S1), describe the feature selection methods considered in this work (Section S2), and present results from our simulation studies (Section S3), cancer studies (Section S4), and sensitivity analyses of the IPSS parameters (Section S5).
S1.  IPSS details
We provide an algorithm that implements integrated path stability selection (IPSS) for feature importance scores (Section S1.1), elaborate on the theory of IPSS and its connection to efp scores (Section S1.2), and describe the IPSS parameters in greater detail (Section S1.3).
S1.1  Algorithm.
Algorithm S1, discussed in Section 2.2, implements IPSS for feature importance scores. The number of grid points used to evaluate the integrals in Algorithm S1 is always $K=100$. Like many of the other IPSS parameters, $K$ is inconsequential provided it is sufficiently large; in our experience, values greater than 25 suffice. This is because the function $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$, the paths $\lambda\mapsto\hat{\pi}_{j}(\lambda)$ (which are monotonically increasing functions of $\lambda$), the quantity $\mathcal{I}(\Lambda)$, and the family of measures $\mu_{\delta}(d\lambda)=z^{-1}_{\delta}\lambda^{-\delta}d\lambda$ are all very numerically stable.
Algorithm S1 Integrated path stability selection for feature importance scores
1:Data $Z_{1:n}$, importance function $\Phi$, number of grid points $K$ and iterations $B$, probability measure $\mu$, function $f$ (default $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$), and cutoff $C$ (default $C=0.05$).
@@
10:while $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])<C$ do
@@
S1.2  IPSS theory and efp scores.
@@
S1.3  Parameters.
Table S1 shows the default IPSS parameters used for IPSSGB and IPSSRF throughout this work (default gradient boosting and random forest parameters are in Section S2.2.1 and Section S2.2.2). As noted above, our choice of function $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ is determined by the availability and strength of the theoretical bound in Theorem S1.1. Similarly, the choice of $\lfloor n/2\rfloor$ samples used to construct the selection probabilities $\hat{\pi}_{j}(\lambda)$ is required for stability selection theorems (not just Theorem S1.1) to hold, and it is unclear how to adapt their proofs to accommodate other sample sizes. Thus, $f$ and $\lfloor n/2\rfloor$ are theoretically determined rather than free parameters.